OpenStudy (anonymous):

one line is a horizontal line. what is the slope of a perpendicular line to it?

OpenStudy (anonymous):

If the slope of the horizontal line = M Then any perpendicular line is equal to = -1/M

OpenStudy (anonymous):

one line has a slope of -3/5. what is the slope of the line perpendicular to it?

OpenStudy (anonymous):

5/3

OpenStudy (anonymous):

find the equation of a horizontal line that has a y-intercep of 5

OpenStudy (anonymous):

A horizontal line that has a point of ( 0 , 5)

OpenStudy (anonymous):

Point - slope form

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First find the Slope of horizontal line

OpenStudy (anonymous):

There's a theorem , That any horizontal line Slope is always equal to zero

OpenStudy (anonymous):

so apply the point slope form Where P(0,5) and m = 0

OpenStudy (anonymous):

y - y1 = m(x-x1)

OpenStudy (anonymous):

0 - y = 0(x - 5)

OpenStudy (anonymous):

y = 0

OpenStudy (anonymous):

wait wrong

OpenStudy (anonymous):

y = 5 is the equation

OpenStudy (anonymous):

vertical lines have an infinite (or undefined) slope

OpenStudy (anonymous):

Yeah But your askin for the equation of a horizontal right